OpenStudy (anonymous):

1.) Two perpendicular lines intersect at the origin. If the slope of the first line is -1/2, what is the equation of the second line? 2.)The slope of a line is -1/2 What is the slope of a line that is parallel to it? 3.)What is the equation of the line that passes through the points (1,2/3) and (-1,-1)

OpenStudy (anonymous):

@diamondprincess @nicolehill @iGreen

OpenStudy (igreen):

Perpendicular lines have opposite slope..what's the opposite of -1/2? @RosePetal_

OpenStudy (sidsiddhartha):

consider two lines \[y=m_x+b\\and\\y=m_2x+c\\they~will~be~perpendicular~when\\\boxed {m_1*m_2=-1}\]

OpenStudy (anonymous):

Wait which number is this for? and opposite of -1/2 is 1/2

OpenStudy (sidsiddhartha):

no perpendicular lines does'nt have only opposite but also negetive slope

OpenStudy (anonymous):

i think i know the answer to number 2 but I'm not completely sure

OpenStudy (igreen):

No, switch 1/2 around what do you get? @RosePetal_

OpenStudy (anonymous):

2

OpenStudy (sidsiddhartha):

so for the first line slope \[m_1=-1/2\] then slope of the line perpendicular to it will be \[m_2=\frac{ -1 }{ m_1 }=\frac{ -1 }{ 1-/2 }=2\] got it? yes good

OpenStudy (igreen):

Yep, so that's the slope of the 2nd line for #1.

OpenStudy (igreen):

For #2: Parallel lines have the same slope..so what's the slope of line 2 gonna be?

OpenStudy (anonymous):

oh okay so for number one you just switch it around

OpenStudy (igreen):

Yep, and you change it to positive..

OpenStudy (anonymous):

wouldn't it be -2

OpenStudy (sidsiddhartha):

now u need the equation for the perpendicular line ok?

OpenStudy (igreen):

2.)The slope of a line is -1/2 What is the slope of a line that is parallel to it? Parallel lines have the SAME slope. So what's the slope of the line parallel to the original? @RosePetal_

OpenStudy (sidsiddhartha):

so equation for the perpendicular line is\[y=2x\]

OpenStudy (anonymous):

2??

OpenStudy (anonymous):

and real fast 2 isn't an answer for number 1 i attached a pic of the answers I'm stuck between a and d

OpenStudy (igreen):

No, 2 is incorrect. Slope of line #1 = -1/2 Slope of line #2 is parallel to line #1. parallel lines have the same slope.

OpenStudy (anonymous):

oh I'm an idiot so it would be -1/2

OpenStudy (igreen):

2 is the slope of the line for #1..and there's only one answer with 2 in it..

OpenStudy (igreen):

Yep, the slope for line #2 will also be -1/2..

OpenStudy (anonymous):

okay sorry I'm not thinking right now a haha and then what about number 3

OpenStudy (sidsiddhartha):

for #3 first calculate slope

OpenStudy (sidsiddhartha):

can u do that?

OpenStudy (sidsiddhartha):

slope=y2-y1/x2-x1

OpenStudy (igreen):

Equation for slope: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) Our points are (1, 0.67) and (-1, -1) Plug them in: \(m = \dfrac{-1-0.67}{-1-1}\) Subtract: \(m = \dfrac{-1.67}{-2}\) Can you divide that? @RosePetal_

OpenStudy (anonymous):

ya it would be .835

OpenStudy (sidsiddhartha):

yes which is equal to 5/6

OpenStudy (anonymous):

so the answer to number 3 is 5/6x-1/6

OpenStudy (igreen):

Yep, so slope is 5/6..

OpenStudy (sidsiddhartha):

now when u have the slope then use \[y-y_1=m(x-x_1)\] this form to calculate equation of the line :)

OpenStudy (anonymous):

another quick question i just wanna double check my answer - Rewrite the slope-intercept form equation into standard form. y = –2x + 4 would the answer be 2x + y = 4

OpenStudy (sidsiddhartha):

yes!

OpenStudy (igreen):

Standard form is Ax + By = C.

OpenStudy (igreen):

So that would be correct. @RosePetal_

OpenStudy (anonymous):

Yas!!! Thanxx so much